OpenStudy (anonymous):

Find the solution to the equation 64^(2 - x) = 4^4x. Using complete sentences, explain the procedure used to solve this equation.

OpenStudy (nurali):

64^(2-x)=4^(4x) 2^(6(2-x))=2^(2(4x)) 6(2-x)=2(4x) 6(2-x)=8x 12-6x=8x 12-14x=0 -14x=-12 x=6/7